A flat surface with area of 7.00 m2 is oriented at an angle q to a uniform electric field with magnitude 6.0 × 105 N/C. If the flux through this surface is 2.0 × 106 N×m2/C, what is q?

81.8°
32.9°
6.18°
61.6°
Answer #1

here,

the area of flat surface , A = 7 m^2

the uniform electric feild , E = 6 * 10^5 N/C

the electric flux , phi = E * A * cos(q)

2 * 10^6 = 6 * 10^5 * 7 * cos(q)

solving for q

q = 61.6 degree

the angle q is 61.6 degree